Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR16 @ 15:30 pm on September 20th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Hall work

  • DC report
  • Water leak status and readiness of conventional magnets
  • Decay solenoid leak status

Operations

Notes

Decay solenoid

There is a leak on the seal - HPC have accepted this and plan to deliver a replacement. The delivery time is 10-12 working days; likely to be fitted in the first or second week of October. Ken Long asks that we push harder, e.g. get them to courier the seal etc, or contact Kaiser directly. MICE will require Linde fridge expert on-site when it is fitted. Once fitted, it will take around 12 days to purify and cool down the fridge. Note that a less resilient seal is also an option, which may arrive more quickly. Also concern that the shaft may be an issue; seek to have a spare. Likely it will take a day to do the seal changeover.

Water leak

Magnets are running at the moment, running at 30 degrees C. Q4 is uncertain. It is likely we can run tomorrow, awaiting confirmation. Richard Hale is investigating earth leakage trip on Q4. It looks like 30 degrees is okay (trying at 200 A). So can get to e.g. 200 MeV/c settings and reference run settings. Nb: Power deposited goes as I^2 R. The issue appears to be that not enough water is going from the roof to the heat exchanger. Need some monitoring of the trench water temperature and alarm, otherwise concerned over potentially terminal heat deposition on the conventional magnets. Note hardware interlock on the conventional magnets (klixons). Pierrick has an input temperature monitor to Q1-3 and Q4-9. Needs to be alarmed. If the alarm sounds and is a temperature alarm, action is for shifters to ramp magnets to 0 and to phone the MOM. Andy Gallagher comments that there are klixons in the magnet and there is a power supply trip, but Henry should check the limits. Power supplies are limited to 35 degrees C on input cooling water.

Henry proposes new alarm limits and requires alarms are in the controls system before shifter magnet operation.

The limitation is that the heat exchanger and the FC cooling are in parallel, so focus coil compressor cooling is tensioned against the conventional magnet cooling. Plan is to take the heat exchanger out of the FC circuit. Plan is to move water for FC to the West wall manifolds; hosing is mostly still in. Need to run a longer water hose from West wall to the mezzanine where the FC compressors are.

Hall plan

Wednesday:-
  • BLOC training 11 am Wednesday morning
  • Water work Wednesday morning
  • Expert-led run (DAQ and conventional beamline) Wednesday afternoon
  • SSD Fan Wednesday evening
Thursday:-
  • Shifter training and water-work Thursday morning
  • Expert-led run (cooling channel magnets) Thursday afternoon

Friday through to Sunday straight track data

Monday onwards:- magnet data
  • Need BLOC for Saturday; Ken Long offers to cover
  • Need shifters from Wednesday onwards.
There was a discussion of the magnet running (for data taking); there will be a meeting Wednesday to shake things down. Questions/needs that arose in discussion...
  • Can the magnets run overnight? some discussion among experts.
  • Establish operational procedures
  • Magnet on-call?
  • Agreed - magnet soak test is next, need to cover the owl shift.

AOB

Diffuser issue - think we need a technician down from Oxford; Victoria to do some digging

Updated by Rogers, Chris about 6 years ago ยท 4 revisions